OpenStudy (anonymous):

In a large bag of marbles, 25% of them are red. A child chooses 4 marbles from this bag. If the child chooses the marbles at random, what is the chance that the child gets exactly three red marbles? A) 0.047 B) 0.141 C) 0.211 D) 0.063 A pet supplier has a stock of parakeets of which 10% are blue parakeets. A pet store orders 3 parakeets from this supplier. If the supplier selects the parakeets at random, what is the chance that the pet store gets exactly one blue parakeet? A) 0.081 B) 0.243 C) 0.027 D) 0.003 In a large bag of marbles, 25% of them are red. A child chooses 4 marbles

OpenStudy (kropot72):

The probability of getting exactly 3 red marbles is found from the binomial distribution. \[P(3\ red)=\left(\begin{matrix}4 \\ 3\end{matrix}\right)0.25^{3} \times 0.75=4 \times 0.25^{3} \times 0.75\]